OpenStudy (anonymous):

HELP!!! :O Which expression describes the rule of this sequence (Question) 7, 14, 21, 28, 35, ? A 7+n B 7n C7^n (Carrying the n ) D n+ 7+7

OpenStudy (anonymous):

b

OpenStudy (anonymous):

7x1 7x2 7x3

OpenStudy (anonymous):

See, here a = 7 and the common difference d = 14 - 7 = 7 So use: \[t_n = a + (n-1).d\] Put the values of a and d here and solve for tn..

OpenStudy (anonymous):

Thanks guys!!!